OpenStudy (anonymous):

What is an equation of the line, in point-slope form, that passes through the given point and has the given slope? point: (8, –8); slope: 3

OpenStudy (anonymous):

@karatechopper

OpenStudy (anonymous):

@help123please.

OpenStudy (anonymous):

@Bevy

OpenStudy (anonymous):

\[\frac{y-y_1}{x-x_1}=m\]

OpenStudy (anonymous):

\[\frac{y+8}{x-8}=3\]

OpenStudy (anonymous):

\[y+8=3(x-8)\]

OpenStudy (anonymous):

not sure but i think -63

OpenStudy (anonymous):

not sure but i think -63

OpenStudy (anonymous):

– 8 = 3 (x + 8) y + 8 = 3 (x + 8) y – 8 = 3 (x – 8) y + 8 = 3 (x – 8)

OpenStudy (anonymous):

@Bevy those are the multiple choice answers. but that you for trying. i dont get an of this at all :p

OpenStudy (anonymous):

this one is the right answer as proven above\[y+8=3(x-8)\]

OpenStudy (anonymous):

take care